• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 22:09:04 on July28,2003

    Entry number 107759

    keyword=NMR: 39.28%

       Moller measurement done, raster checked with spot++
    and adjusted.

    21:52    Move to 3He cell position.

    21:58    NMR measurement done, lasers on, but
    we are in the 3He cell position, not the pickup coils
    position (see HAlog entry


    Result: polarization = 39.28%

    Leave target as/is, ask for beam back.

    (aside: lots of thunder heard outside, let's hope the beam stays on!)